# Sparsity-Promoting Dynamic Mode Decomposition Applied to Sea Surface Temperature Fields

### v1

@@ I Introduction
In recent decades, the operator-theoretic approach, with a focus on the Koopman operator [5], has gained great attention for analyzing dynamical systems by lifting nonlinear dynamics into a linear representation. This state-of-the-art method describes the linear evolution of observables in an infinite-dimensional functional space [6, 7]. Needless to say, applied Koopmanism played a growing role in uncovering structures in (geophysical) fluid dynamics [8, 9] and in interpreting coherent modes in climate and weather systems. When Koopman mode decomposition (KMD) [10] is combined with its numerical counterparts, such as various dynamic mode decomposition (DMD) [11, 12, 13], it provides spectral analysis for decomposing spatiotemporal patterns. Related literature has validated these methods for capturing climate-related phenomena, including the Madden–Julian Oscillation (MJO) [14] and its dominant component, tropical intraseasonal variability [15, 16], the El Niño–Southern Oscillation (ENSO) [17, 18], sea surface temperature (SST) anomalies [19], the Pacific Decadal Oscillation (PDO) [20], and sea ice cover variability [21]. Notably, Navarra et al. [19] estimated the Perron-Frobenius and Koopman operators based on one-dimensional Niño-3 time series data and monthly mean Pacific SST, where the system evolution density was derived from the transfer or Koopman eigenfunctions. Subsequently, the variability of tropical ocean and PDO of global SST were identified via Koopman modes [20], and extended to examine the continuous spectrum [6, 22].
@@
The rest of this paper is organized as follows: Section II introduces the DMD method for approximating the KMD, and deriving a low-dimensional linear system via SPDMD. Section III presents three experiments on SST data, extracting dominant Koopman modes to analyze spatial and temporal structures in climate systems. Section IV concludes the paper. Finally, the Koopman operator and other details for continuous spectra are given in Appendix.
@@ II-B Sparsity-Promoting Dynamic Mode Decomposition
After solving the sparse optimization problem (7) or (6), a sparse solution for the amplitudes ${\bf{b}}^{\star}$ is obtained, with cardinality $\mathrm{card}({\bf{b}}^{\star}(\gamma))=s$ (or $\ell_{0}$ norm $\|{\bf{b}}^{\star}(\gamma)\|_{0}=s$). This cardinality indicates the number of non-zero amplitudes corresponding to the most dominant modes, ordered by their magnitudes $|b_{i}|$ for $i=1,\ldots,s$. Letting $m\triangleq\min\{s,r\}$, where $r$ is the SVD truncation rank in (4), we obtain a reduced-order model using only the selected $m$ modes.
@@
II-C Model-Reduction: A Low Dimensional Dynamics
To realize a low-dimensional representation of the original discrete-time system (19), we project the dynamics onto a subspace spanned by a selected set of $m$ Koopman modes (e.g., see [10]) using sparsity-based selection, expressed as
${\bf{g}}({\bf{x}}_{k})\approx\sum_{j=1}^{m}\lambda_{j}^{t}\varphi_{j}({\bf{x}}_{0}){\bf{v}}_{j}=\sum_{j=1}^{m}\eta_{j}[k]{\bf{v}}_{j}:={\bf{V}}{\bm{\eta}}[k],$ (9)
where ${\bm{\eta}}=[\eta_{1},\ldots,\eta_{m}]^{\top}\in{{\mathbb{C}}}^{m}$ and ${\bf{V}}=[{\bf{v}}_{1},\ldots,{\bf{v}}_{m}]\in{\mathbb{C}}^{p\times{m}}$. By substituting (9) into (21), we have
$\sum_{j=1}^{m}\eta_{j}[k+1]{\bf{v}}_{j}=\big{(}{\bf{g}}\circ{{\bf{T}}}\circ{\bf{g}}^{-1}\big{)}({\bf{V}}{\bm{\eta}}[k]),$ (10)
where the inverse of observable function ${\bf{g}}^{-1}$ exists because the data composed of observables is always available through direct measurements in climate systems, such as SST data.
By multiplying both sides of (10) by ${\bf{v}}_{i}^{{\dagger}}$, we have
$\sum_{j=1}^{m}\eta_{j}[k+1]{\bf{v}}_{i}^{{\dagger}}{\bf{v}}_{j}$ $=\underbrace{{\bf{v}}_{i}^{{\dagger}}\big{(}{\bf{g}}\circ{{\bf{T}}}\circ{\bf{g}}^{-1}\big{)}({\bf{V}}{\bm{\eta}}[k])}_{:=R_{i}({\bm{\eta}}[k])}$ (11)
for $i=1,2,\ldots,m$. Now if we take $\widehat{{\bf{V}}}:={\bf{V}}^{{\dagger}}{\bf{V}}\in{\mathbb{C}}^{m\times{m}}$, it thus gives rise to a compact form of a low-dimensional linear system representation of (11), that is,
$\widehat{{\bf{V}}}\eta[k+1]={\bf{R}}({\bm{\eta}}[k]),$ (12)
where ${\bf{R}}({\bm{\eta}}):=[R_{1}({\bm{\eta}}),\ldots,R_{m}({\bm{\eta}})]^{\top}\in{\mathbb{C}}^{m}$. This discrete time system (11) or (12) defined on ${\mathbb{C}}^{m}$ gives a reduced-order system model of (19) in terms of the first $m$ Koopman modes, which can be achieved through SPDMD (7).
@@ III-A Data Source: Long-Term SST Data
In this context, the (long-term) time-series dataset are obtained from monthly SST over Niño-3 index (https://psl.noaa.gov/gcos_wgsp/Timeseries/Nino3), taking only sea mask into account (without land data information), see in Fig. 2; For more details, we refer the reader to the available data source and the related animation illustrating the time evolution of the SST data [25].
@@ III-D Experiment 3: Annual Cycle of SST Data Field
Figure 8: The magniude of the spatial modes for annual cycle of SST.
@@ IV-B Koopman Mode Decomposition
where ${\bf{x}}_{k}={\bf{x}}(kh)={\bf{x}}(t_{k})$ and the discrete-time map $S^{t}_{h}(\cdot):={\bf{T}}(\cdot)$ evolves ${\bf{x}}_{k}$ to ${\bf{x}}_{k+1}$ [10]. Analogous expression holds for discrete-time Koopman eigenvalue (KE) and Koopman eigenfunction (KEF), (18) replaced by ${\mathcal{K}}{\varphi}_{j}={\lambda}_{j}{\varphi}_{j}$ if we assume ${\lambda}_{j}=e^{\nu_{j}{t}}$. Moreover, the vector-valued observable ${\bf{g}}:{\mathcal{M}}\to{\mathbb{R}}^{m}$ defined by
${\bf{g}}({\bf{x}})=[g_{1}({\bf{x}}),\cdots,g_{m}({\bf{x}})]^{\top},$ (20)
can be expanded in terms of diverse combinations of the eigenfunctions $\{\varphi_{j}({\bf{x}})\}_{j=1}^{\infty}$. In this context, the vector-valued full-state observables ${\bf{y}}={\bf{g}}({\bf{x}})={\bf{x}}$ when $p=n$, simply returns the state vector itself with the same dimensionality. For example, if all the elements of observable ${\bf{g}}$ in (20) lie within the span of the eigenfunctions $\varphi_{j}$, then it can be derived in terms of a (countable) set of these KEFs ${\bf{g}}({\bf{x}})=\sum_{j=1}^{\infty}\varphi_{j}({\bf{x}}){\bf{v}}_{j}$, where ${\bf{v}}_{j}\in{\mathbb{C}}^{p}$ is the well-known Koopman modes (KM) as a vector-valued coefficient for the expansion. Besides, Koopman operator $\mathcal{K}$ evolves the observable according to (17), yielding the Koopman mode decomposition (KMD) as follows
@@
where the elements $\{\lambda_{j},\varphi_{j},{\bf{v}}_{j}\}$, $j=1,2,\ldots,$ or $\infty$ are referred to as Koopman tuples. The Koopman eigenfunctions and Koopman eigenvalues encode lots of information about the underlying dynamical systems. Finite-dimensional approximations of the Koopman operator $\mathcal{K}$ rely on a finite number of tuples as defined in (21), which can be computed using projection-based, data-driven DMD methods (see (5) and [11] for more details).
Figure 9: Pseudospectra computed for the seasonal cycle of SST data. The blue points represent the kernelized EDMD eigenvalues computed using the Laplacian kernel (24).
IV-C Another Look at Continuous Spectrum
Suppose that there is no singular continuous spectrum. Then any (scalar) observable $g\in{\mathcal{L}}^{2}(\mathcal{M},\mu)$, it follows that
$g=\sum_{\lambda_{j}\in\sigma(\mathcal{K})}\varphi_{j}{\bf{v}}_{j}+\int_{[-\pi,\pi]_{\text{per}}}\phi_{\theta,g}{\mathrm{d}}\theta$ (22)
where $\sigma(\lambda)$ is the set of eigenvalues (i.e., spectrum) of $\mathcal{K}$, and $\phi_{\theta,g}$ is a “continuously parameterized” collection of eigenfunctions [22, Sec. 2]. Based on (21), the KMD can be further defined as [6]
$g({\bf{x}}_{t})$ $={\mathcal{K}}^{t}(g({\bf{x}}_{0}))$
$=\mathop{\sum}\limits_{\lambda_{j}\in\sigma(\mathcal{K})}\lambda_{j}^{t}\varphi_{j}({\bf{x}}_{0}){\bf{v}}_{j}+\int_{[-\pi,\pi]_{\text{per}}}e^{in\theta}\phi_{\theta,g}({\bf{x}}_{0}){\mathrm{d}}\theta.$
Given a dictionary $\{\psi_{1},\psi_{2},\ldots,\psi_{N_{k}}\}$ of observables, the extended dynamic mode decomposition (EDMD) [12] using Galerkin method selects a matrix ${\bf{A}}_{\rm{EDMD}}$ that approximates $\mathcal{K}$ on the subspace ${\mathcal{F}}_{N_{k}}:={\rm{span}}\{\psi_{1},\psi_{2},\ldots,\psi_{N_{k}}\}$, that is,
${\mathcal{K}}\psi_{j}({\bf{x}})=\psi({\bf{T}}({\bf{x}}))\approx\sum_{i=1}^{N_{k}}({\bf{A}}_{\rm{EDMD}})_{ij}\psi_{i}({\bf{x}})$
for $1\leq{j}\leq{N}_{k}$. Define the vector-valued feature map $\Psi({\bf{x}})=[\psi_{1}({\bf{x}}),\ldots,\psi_{N_{k}}({\bf{x}})]\in{\mathbb{C}}^{1\times{N_{k}}}$. Then for any observable $g\in{\mathcal{F}}_{N_{k}}$ can be written as $g({\bf{x}})=\sum_{j=1}^{N_{k}}\psi_{j}({\bf{x}})g_{j}=\Psi({\bf{x}}){{\bf{g}}}$ for some vector ${{\bf{g}}}=[g_{1},\ldots,g_{N_{K}}]^{\top}\in{\mathbb{C}}^{N_{k}}$. We have
$\quad\mathcal{K}g({\bf{x}})=\Psi({\bf{T}}({\bf{x}})){\bf{g}}$
$=\Psi({\bf{A}}_{\rm{EDMD}}{\bf{g}})\!+\!\Bigg{(}\sum_{j=1}^{N_{k}}\psi_{j}({\bf{T}}({\bf{x}}))g_{j}-\Psi({\bf{x}})({\bf{A}}_{\rm{EDMD}}{\bf{g}})\Bigg{)}.$
Furthermore, kernelized EDMD [13] applies the kernel trick for EDMD setting to compute ${\bf{A}}_{\rm{EDMD}}$ by introducing a positive-definite kernel function $\kappa:\mathcal{M}\times\mathcal{M}\to{\mathbb{R}}$, which induces a feature map ϝ$\digamma:{\mathbb{R}}^{n}\to\mathcal{H}$ so that
ϝ$\langle\digamma({\bf{x}}),\digamma({\bf{x}}^{\prime})\rangle=\kappa({\bf{x}},{\bf{x}}^{\prime}),$
where $\mathcal{H}$ represents the reproducing kernel Hilbert space.
Figure 10: Pseudospectra computed for the annual cycle of SST data. The shaded bone colormap indicates the threshold values $\epsilon$. The red points represent the kernelized EDMD eigenvalues computed using the Gaussian kernel (23).
It is clear that different choices of kernel result in different matrices ${\bf{A}}_{\rm EDMD}$ and in general different eigenvectors and eigenfunctions. In this context, we consider the following (nonlinear) kernel function cases for a suitable parameter $c$:
  • 1)
    Gaussian kernel:
    $\kappa({\bf{x}},{\bf{x}}^{\prime})=\exp(-\|{\bf{x}}-{\bf{x}}^{\prime}\|^{2}/c^{2}).$ (23)
  • 2)
    Laplacian kernel:
    $\kappa({\bf{x}},{\bf{x}}^{\prime})=\exp(-\|{\bf{x}}-{\bf{x}}^{\prime}\|/c).$ (24)
For a finite matrix ${\bf{A}}$ and $\epsilon>0$, the $\epsilon$-pseudospectrum of ${\bf{A}}$ is defined as
$\sigma_{\epsilon}({\bf{A}})$ =$=\left\{\lambda\in{\mathbb{C}}:\|({\bf{A}}-\lambda{I})^{-1}\|\geq{1}/\epsilon\right\}$
$=\cup_{\|{\bf{B}}\|\leq\epsilon}\sigma({\bf{A}}+{\bf{B}}).$
It should be noted that if we replace ${\bf{A}}$ with the Koopman operator $\mathcal{K}$, we need consider the closure of the set for $\sigma_{\epsilon}(\mathcal{K})$, as the Koopman operator may be an unbounded operator.
Given the grids $z_{1},z_{2},\ldots,z_{k}\in{\mathbb{C}}$, one can approximate $\epsilon$-pseudospectrum $\{z_{j}:\tau_{j}<\epsilon\}$ and approximiate the eigenfunction $\{\varpi_{j}:\tau_{j}<\epsilon\}$ such that ${\bf{A}}\varpi_{j}=\lambda_{j}\varpi_{j}$, where $\tau_{j}$ can be viewed as the resudials. As mentioned in [22], residual DMD is a powerful tool for computing the pseudo-spectrum and continuous spectrum based on EDMD and kernelized EDMD via the Galerkin method; see also the references therein for detailed algorithms.
By employing residual DMD, we perform residual DMD on the SST data fields. Fig. 9 shows the resudials and pesudospestrum of the seasonal cycle of the SST data field, where the residuals (blue points) are plotted against the theoretical curve $\sqrt{1-|\lambda|^{2}}$ (black), corresponding to the regression under the (nonlinear) Laplacian kernel function. Similarity, Fig. 10 displays the resudials and pesudospestrum of the annual cycle of SST data using the Gaussian kernel, while the residuals (red points) lie far from th curve $\sqrt{1-|\lambda|^{2}}$.

### v2

@@ I Introduction
In recent decades, the operator-theoretic approach, with a focus on the Koopman operator [5], has gained great attention for analyzing dynamical systems by lifting nonlinear dynamics into a linear representation. This state-of-the-art method describes the linear evolution of observables in an infinite-dimensional functional space [6, 7]. Needless to say, applied Koopmanism played a growing role in uncovering structures in (geophysical) fluid dynamics [8, 9] and in interpreting coherent modes in climate and weather systems. When Koopman mode decomposition (KMD) [10] is combined with its numerical counterparts, such as various dynamic mode decomposition (DMD) [11, 12], it provides spectral analysis for decomposing spatiotemporal patterns. Related literature has validated these methods for capturing climate-related phenomena, including the Madden–Julian Oscillation (MJO) [13] and its dominant component, tropical intraseasonal variability [14, 15], the El Niño–Southern Oscillation (ENSO) [16, 17], sea surface temperature (SST) anomalies [18], the Pacific Decadal Oscillation (PDO) [19], and sea ice cover variability [20]. Notably, Navarra et al. [18] estimated the Perron-Frobenius and Koopman operators based on one-dimensional Niño-3 time series data and monthly mean Pacific SST, where the system evolution density was derived from the transfer or Koopman eigenfunctions. Subsequently, the variability of tropical ocean and PDO of global SST were identified via Koopman modes [19], and extended to examine the continuous spectrum [6, 21].
@@
The rest of this paper is organized as follows: Section II introduces the DMD method for approximating the KMD, and deriving a low-dimensional linear system via SPDMD. Section III presents three experiments on SST data, extracting dominant Koopman modes to analyze spatial and temporal structures in climate systems. Section IV concludes the paper. Finally, some preliminaries on the Koopman operator are provided in the Appendix for interested readers.
@@ II-B Sparsity-Promoting Dynamic Mode Decomposition
After solving the sparse optimization problem (7) or (6), a sparse solution for the amplitudes ${\bf{b}}^{\star}$ is obtained, with cardinality $\mathrm{card}({\bf{b}}^{\star}(\gamma))=s$ (or $\ell_{0}$ norm $\|{\bf{b}}^{\star}(\gamma)\|_{0}=s$). This cardinality indicates the number of non-zero amplitudes corresponding to the most dominant modes, ordered by their magnitudes $|b_{i}|$ for $i=1,\ldots,s$. Letting $m\triangleq\min\{s,r\}$, where $r$ is the SVD truncation rank in (4), we can consider a reduced-order model using only the selected $m$ modes.
@@ III-A Data Source: Long-Term SST Data
In this context, the (long-term) time-series dataset are obtained from monthly SST over Niño-3 index [^1] https://psl.noaa.gov/gcos_wgsp/Timeseries/Nino3, taking only sea mask into account (without land data information), see in Fig. 2; For more details, we refer the reader to the available data source and the related animation illustrating the time evolution of the SST data [^2] SST data and animation are available..
@@ III-D Experiment 3: Annual Cycle of SST Data Field
Figure 8: The magnitude of the spatial modes for annual cycle of SST.
@@ IV-B Koopman Mode Decomposition
where ${\bf{x}}_{k}={\bf{x}}(kh)={\bf{x}}(t_{k})$ and the discrete-time map $S^{t}_{h}(\cdot):={\bf{T}}(\cdot)$ evolves ${\bf{x}}_{k}$ to ${\bf{x}}_{k+1}$ [10]. Analogous expression holds for discrete-time Koopman eigenvalue (KE) and Koopman eigenfunction (KEF), (14) replaced by ${\mathcal{K}}{\varphi}_{j}={\lambda}_{j}{\varphi}_{j}$ if we assume ${\lambda}_{j}=e^{\nu_{j}{t}}$. Moreover, the vector-valued observable ${\bf{g}}:{\mathcal{M}}\to{\mathbb{R}}^{p}$ defined by
${\bf{g}}({\bf{x}})=[g_{1}({\bf{x}}),\cdots,g_{p}({\bf{x}})]^{\top},$ (16)
can be expanded in terms of diverse combinations of the eigenfunctions $\{\varphi_{j}({\bf{x}})\}_{j=1}^{\infty}$. In this context, the vector-valued full-state observables ${\bf{y}}={\bf{g}}({\bf{x}})={\bf{x}}$ when $p=n$, simply returns the state vector itself with the same dimensionality. For example, if all the elements of observable ${\bf{g}}$ in (16) lie within the span of the eigenfunctions $\varphi_{j}$, then it can be derived in terms of a (countable) set of these KEFs ${\bf{g}}({\bf{x}})=\sum_{j=1}^{\infty}\varphi_{j}({\bf{x}}){\bf{v}}_{j}$, where ${\bf{v}}_{j}\in{\mathbb{C}}^{p}$ is the well-known Koopman modes (KM) as a vector-valued coefficient for the expansion. Besides, Koopman operator $\mathcal{K}$ evolves the observable according to (13), yielding the Koopman mode decomposition (KMD) [6, 7] as follows
@@
where the elements $\{\lambda_{j},\varphi_{j},{\bf{v}}_{j}\}$, $j=1,2,\ldots,$ or $\infty$ are referred to as Koopman tuples. The Koopman eigenfunctions and Koopman eigenvalues encode lots of information about the underlying dynamical systems. Finite-dimensional approximations of the Koopman operator $\mathcal{K}$ rely on a finite number of tuples as defined in (17), which can be computed using projection-based, data-driven DMD methods (see (5) and [3, 11] for more details).